# Explaining Tournament Solutions with Minimal Supports

### v2

@@ 11          $\mu_{\mathcal{X}}(c^{\prime},c)\leftarrow n-t$
A.3 Definitions and Proofs for Section 4.4
We start this section by presenting useful results characterising necessary winners in terms of the score margins of the winning candidate, a notion that will be used to construct our algorithms for the computation of minimal supports. Our results are for weighted tournaments and the Borda rule, but they also apply to Copeland by setting $n=1$.
@@
Then, we can introduce the notion of score margin, the main notion that we will used in the proofs of this paper.
@@
We can then provide a characterisation for necessary winner in terms of score margin, in line with an analogous result by Konczak and Lang (2005) in voting.
@@
Let $\mathcal{X}$ be a SMS for $w\in\operatorname{BO}(G)$, by definition $\mathcal{X}$ is a MS for $w\in\operatorname{BO}(G)$. Summing the inequalities from Proposition 23 we have:
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$. For all MSs $\mathcal{X}$ for $w\in\operatorname{BO}(G)$, we have $|\mathcal{X}|\geq n(m-1)-\lfloor n(1-1/m)\rfloor$.
@@
Let $\mathcal{X}$ be a MS for $w\in\operatorname{BO}(G)$. For $w$ to be a necessary winner, according to Proposition 23, we need $\forall c\in\mathcal{C}\setminus{\{w\}},\,\sigma^{min}_{\operatorname{BO}}(w,\mathcal{X})\geq\sigma^{max}_{\operatorname{BO}}(c,\mathcal{X})$. This gives us the constraint $\sigma^{min}_{\operatorname{BO}}(w,\mathcal{X})\geq\mathbb{E}_{c\neq w}\sigma^{max}_{\operatorname{BO}}(c,\mathcal{X})$. Since there always exists a losing candidate with score at least as high as the mean of the losing candidates, the winner’s score must be larger than the average of the losers’ scores. From this inequality we can derive a second lower bound on the size of a SMS.
@@
To conclude, note that $\sum_{c\neq w}\mu_{\mathcal{X}}(w,c)\leq\sum_{c^{\prime}\in\mathcal{C}}\sum_{c\in\mathcal{C}}\mu_{\mathcal{X}}(c^{\prime},c)=|\mathcal{X}|$. ∎
@@
Let $\mathcal{X}$ be a SMS for $w\in\operatorname{BO}(G)$. For $w$ to be a necessary winner, according to Proposition 23, we need $\forall c\in\mathcal{C}\setminus{\{w\}},\,\sigma^{min}_{\operatorname{BO}}(w,\mathcal{X})\geq\sigma^{max}_{\operatorname{BO}}(c,\mathcal{X})$. We have m1m-1 local constraints (one for each losing candidate) which can be combined as $\sigma^{min}_{\operatorname{BO}}(w,\mathcal{X})\geq\max_{c\neq w}\sigma^{max}_{\operatorname{BO}}(c,\mathcal{X})$. This transforms into $|\mathcal{X}|\geq\sum_{c}\mu_{\mathcal{X}}(w,c)\geq n(m-1)-\min_{c\neq w}(\mu_{\mathcal{X}}(w,c))$. Finally, since $\mu_{\mathcal{X}}\leq\mu$, $|\mathcal{X}|\geq n(m-1)-\min_{c\neq w}(\mu(w,c))$. ∎
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$ with a Borda score of $\sigma_{w}$. If $\sigma_{w}\leq n(m-1)-\min_{c\neq w}\mu(w,c)$, there exists a MS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $|\mathcal{X}|{=}(m-1)(n(m-1)-\sigma_{w})$.
@@
Suppose $\sigma_{w}\leq n(m-1)-\min_{c\neq w}\mu(w,c)$. Hence, for all $c\in\mathcal{C}\setminus{\{w\}}$, $\delta(c,{E}^{w\rightarrow})=\sigma_{\operatorname{BO}}^{min}(c,G)-\sigma_{\operatorname{BO}}^{max}(c^{\prime},G)=\sigma_{w}-(n(m-1)-\mu(w,c))\leq 0$ and ${E}^{w\rightarrow}$ is at best a MS if all deltas are null else it is a subset of a MS. We start with $\mathcal{X}={E}^{w\rightarrow}$. The idea is then to simply add defeats to candidates that can still achieve more victories than the tournament winner, i.e., for each $c\in\mathcal{C}\setminus{\{w\}}$, add elements of ${E}^{\rightarrow c}$ until $\delta(c,{E}^{w\rightarrow})=0$. In the resulting minimal support $\mathcal{X}$, each losing candidate would have $(n(m-1)-\sigma_{w})$ defeats among which all the tournament winner victories can be found. Thus, $|\mathcal{X}|{=}(m-1)(n(m-1)-\sigma_{w})$. ∎
Before working we the other cases, we show the existence of a SMS which is either included in ${E}^{w\rightarrow}$ or contains it.
@@
Given $n$ voters, an $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$, there exists a SMS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $\mathcal{X}\subseteq{E}^{w\rightarrow}$ or ${E}^{w\rightarrow}\subseteq\mathcal{X}$.
@@
The idea is to go from one SMS to another by adding a win in pairwise comparison for $w$ and removing a lose for another candidate. This can be done as long as there are win for $w$ not in the SMS and defeats for other candidates in the SMS. When one of those two conditions is not met anymore we stop. First we show that: (a) For all SMS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $\exists e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\in E\setminus{{E}^{w\rightarrow}}$ and $\exists e_{w}\in{E}^{w\rightarrow}$ with $e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\in\mathcal{X}$ and $e_{w}\not\in\mathcal{X}$, $\left(\mathcal{X}\setminus{\{e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\}}\right)\cup\{e_{w}\}$ is another SMS for $w\in\operatorname{BO}(G)$. Let $\mathcal{X}^{\prime}=\left(\mathcal{X}\setminus{\{e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\}}\right)\cup\{e_{w}\}$. Let $c_{0}\in\mathcal{C}$ s.t. $e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\in{E}^{c_{0}\rightarrow}$. We have $\sigma^{min}(c,\mathcal{X}^{\prime})=\sigma^{min}(c,\mathcal{X})+1$, for all $c\in\mathcal{C}\setminus{\{w,c_{0}\}}$, $\sigma^{max}(c,\mathcal{X})=\sigma^{max}(c,\mathcal{X}^{\prime})$ and $\sigma^{max}(c_{0},\mathcal{X})=\sigma^{max}(c,\mathcal{X}^{\prime})+1$. All in all, for all $c\in\mathcal{C}\setminus{\{w\}}$, $\delta(c,\mathcal{X}^{\prime})\geq\delta(c,\mathcal{X})$. Hence, $\mathcal{X}^{\prime}$ is a weak MS for $w\in\operatorname{BO}(G)$. Additionally, $|\mathcal{X}|=|\mathcal{X}^{\prime}|$ and $\mathcal{X}$ is a SMS for $w\in\operatorname{BO}(G)$. Thus, $\mathcal{X}^{\prime}$ is one too.
Finally, given any SMS $\mathcal{X}$ of $w\in\operatorname{BO}(G)$, applying (a) $\min(|\mathcal{X}\setminus{{E}^{w\rightarrow}}|,|{E}^{w\rightarrow}\setminus{\mathcal{X}}|)$ times, we obtain a SMS for $w\in\operatorname{BO}(G)$ such that $\mathcal{X}\subseteq{E}^{w\rightarrow}$ or ${E}^{w\rightarrow}\subseteq\mathcal{X}$. ∎
@@
Let $\mathcal{X}$ be a SMS for $w\in\operatorname{BO}(G)$. Every element of ${E}^{w\rightarrow}$ contributes 1 to every delta (as a victory for the tournament winner, it increases its minimal possible score by 1) and 1 to the specific candidate $c$ it is pointing toward ((w,c) as defeat for c decreases its maximal possible score by 1). Hence, removing an element of ${E}^{w\rightarrow}$ reduces one delta by 2 and the others by 1. By subset minimality its is not possible to remove any extra elements from $\mathcal{X}$ either when there is no delta greater than 2 or a null delta. ∎
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$ with a Borda score of $\sigma_{w}$. If $\sigma_{w}>n(m-1)-\min_{c\neq w}\mu(w,c)$ and $\forall c\in\mathcal{C}\setminus{\{w\}},\,\delta(c,\mathcal{X})\leq 1$, there exists a MS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $|\mathcal{X}|=n(m-1)-\lfloor n(1-1/m)\rfloor$.
@@
Lemma 28 gives us the existence of a SMS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $\mathcal{X}\subseteq{E}^{w\rightarrow}$. Suppose $\forall c\in\mathcal{C}\setminus{\{w\}},\,\delta(c,\mathcal{X})\leq 1$.
@@
Observe that the left hand-side of the last equality is an integer and since (m1)/m(m-1)/m is strictly less than 1 and non zero ($m>2$), $\sum_{c\neq w}\mu_{\mathcal{X}}(w,c)\leq n(m-1)^{2}/m+(m-1)/m\implies|\mathcal{X}|=\sum_{c\neq w}\mu_{\mathcal{X}}(w,c)\leq\lceil n(m-1)^{2}/m\rceil$. ∎
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$ with a Borda score of $\sigma_{w}$. If $\sigma_{w}>n(m-1)-\min_{c\neq w}\mu(w,c)$ and $\exists c_{0}\in\mathcal{C}\setminus{\{w\}},\,\delta(c_{0},\mathcal{X})\leq 0$, there exists a MS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $|\mathcal{X}|\leq n(m-1)-\min_{c\neq w}\mu(w,c)$.
@@
Lemma 28 gives us the existence of a SMS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $\mathcal{X}\subseteq{E}^{w\rightarrow}$. Suppose $\exists c_{0}\in\mathcal{C}\setminus{\{w\}},\,\delta(c_{0},\mathcal{X})\leq 0$. Since $\mathcal{X}$ is a MS, $\forall c\in\mathcal{C}\setminus{\{w\}},\,\delta(c,\mathcal{X})\geq 0$. Hence, $\delta(c_{0},\mathcal{X})=0$, i.e., $\sum_{c\in\mathcal{C}\setminus{\{w\}}}\mu_{\mathcal{X}}(w,c)=n(m-1)-\mu_{\mathcal{X}}(w,c_{0})$. Additionally, $\forall c\in\mathcal{C}\setminus{\{w\}},\,\delta(c,\mathcal{X})\geq 0=\delta(c_{0},\mathcal{X})$ and we have $\forall c\in\mathcal{C}\setminus{\{w\}},\,\mu_{\mathcal{X}}(w,c)\geq\mu_{\mathcal{X}}(w,c_{0})$. Hence, we have $\forall c\in\mathcal{C}\setminus{\{w\}},\,|\mathcal{X}|\leq n(m-1)-\mu_{\mathcal{X}}(w,c)$ and in particular, $|\mathcal{X}|\leq n(m-1)-\min_{c\in\mathcal{C}\setminus{\{w\}}}\mu_{\mathcal{X}}(w,c)$.
@@
Note that in the last formula we have $\min_{c\in\mathcal{C}\setminus{\{w\}}}\mu_{\mathcal{X}}(w,c)$ and not $\min_{c\in\mathcal{C}\setminus{\{w\}}}\mu(w,c)$. We still have to show that there is a SMS such that there exists a candidate $c$ distinct from $w$ such that $\mu_{\mathcal{X}}(w,c)=\mu(w,c)$.
@@
Suppose for all SMS $\mathcal{Y}\subseteq{E}^{w\rightarrow}$, $|\mathcal{Y}|>n(m-1)-\min(\lfloor n(1-1/m)\rfloor,\min_{c\neq w}\mu(w,c))$. If $\forall c\in\mathcal{C}\setminus{\{w\}}$, $\delta(c,\mathcal{Y})\leq 1$, we are done thanks to Lemma 30. Otherwise, we have the existence of $c_{1}\neq w$ such that $\delta(c_{1},\mathcal{Y})>1$ and by minimality of MSs, $\exists c_{0}\neq w$ such that $\delta(c_{0},\mathcal{Y})=0$. Suppose now that $\forall c\neq w$ such that $\delta(c,w)=0$, $\mu_{\mathcal{Y}}(w,c)<\mu(w,c)$. Consider $\mathcal{Y}^{\prime}$ such that $\mu_{\mathcal{Y}^{\prime}}(w,c_{0})=\mu_{\mathcal{Y}}(w,c_{0})+1$, $\mu_{\mathcal{Y}^{\prime}}(w,c_{1})=\mu_{\mathcal{Y}}(w,c_{1})-1$ (which is possible as $\mu_{\mathcal{Y}^{\prime}}(w,c_{1})<\mu(w,c_{1})$) and else $\mu_{\mathcal{Y}^{\prime}}=\mu_{\mathcal{Y}}$. $\delta(c_{0},\mathcal{Y}^{\prime})=1$, $\delta(c_{1},\mathcal{Y}^{\prime})=\delta(c_{1},\mathcal{Y})-1\geq 1$ and $\forall c\not\in\{w,c_{0},c_{1}\},\,\delta(c,\mathcal{Y}^{\prime})=\delta(c_{1},\mathcal{Y})\geq 1$. Since all deltas are positive, $\mathcal{Y}^{\prime}$ is a weak minimal support and since $|\mathcal{Y}^{\prime}|=|\mathcal{Y}|$ and $\mathcal{Y}$ is a SMS, $\mathcal{Y}^{\prime}$ is a SMS too. By passing from $\mathcal{Y}$ to $\mathcal{Y}^{\prime}$ we strictly reduce the number of losing candidates with a null delta ($\delta(c_{0},\mathcal{Y})=0$ but $\delta(c_{0},\mathcal{Y}^{\prime})=1$). Since the number of candidates is finite, this process has to stop at one point, i.e., $\exists\mathcal{Y}^{*}$ a SMS such that $\exists c^{*}\neq w$ with $\delta(c^{*},\mathcal{Y}^{*})=0$ and $\mu_{\mathcal{Y}}(w,c^{*})=\mu(w,c^{*})$. We have $\sum_{c\neq w}\mu_{\mathcal{Y}^{*}}(w,c^{*})-n(m-1)+\mu(w,c^{*})=0$ and $|\mathcal{Y}^{*}|=n(m-1)-\mu(w,c^{*})$. Finally, since $\min_{c\neq w}\mu(w,c)\leq\mu(w,c^{*})$, exists a SMS $\mathcal{X}$ such that, $|\mathcal{X}|\leq n(m-1)-\min_{c\neq w}\mu(w,c)$. ∎

### v3

@@ 11          $\mu_{\mathcal{X}}(c^{\prime},c)\leftarrow n-t$
A.4 Definitions and Proofs for Section 4.4
We start this section by presenting useful results characterizing necessary winners in terms of the score margins of the winning candidate, a notion that will be used to construct our algorithms for the computation of minimal supports. Our results are for weighted tournaments and the Borda rule, but they also apply to Copeland by setting $n=1$.
@@
Then, we can introduce the notion of score margin, the main notion that we will used in the proofs of this section.
@@
We can then provide a characterisation for necessary winner in terms of score margin, in line with an analogous result by konczak_voting_2005 in voting.
@@
Let $\mathcal{X}$ be an MS for $w\in\operatorname{BO}(G)$, by definition $\mathcal{X}$ is an MS for $w\in\operatorname{BO}(G)$. Summing the inequalities from Proposition 23 we have:
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$. For all MSs $\mathcal{X}$ for $w\in\operatorname{BO}(G)$, we have $|\mathcal{X}|\geq\lceil n(m-1)^{2}/m\rceil$.
@@
Let $\mathcal{X}$ be an MS for $w\in\operatorname{BO}(G)$. For $w$ to be a necessary winner, according to Proposition 23, we need $\forall c\in\mathcal{C}\setminus{\{w\}},\,\sigma^{min}_{\operatorname{BO}}(w,\mathcal{X})\geq\sigma^{max}_{\operatorname{BO}}(c,\mathcal{X})$. This gives us the constraint $\sigma^{min}_{\operatorname{BO}}(w,\mathcal{X})\geq\mathbb{E}_{c\neq w}\sigma^{max}_{\operatorname{BO}}(c,\mathcal{X})$. Since there always exists a losing candidate with score at least as high as the mean of the losing candidates, the winner’s score must be larger than the average of the losers’ scores. From this inequality we can derive a second lower bound on the size of an SMS.
@@
To conclude, note that
$\sum_{c\neq w}\mu_{\mathcal{X}}(w,c)\leq\sum_{c^{\prime}\in\mathcal{C}}\sum_{c\in\mathcal{C}}\mu_{\mathcal{X}}(c^{\prime},c)=|\mathcal{X}|$. ∎
@@
Let $\mathcal{X}$ be an MS for $w\in\operatorname{BO}(G)$. For $w$ to be a necessary winner, according to Proposition 23, we need $\forall c\in\mathcal{C}\setminus{\{w\}},\,\sigma^{min}_{\operatorname{BO}}(w,\mathcal{X})\geq\sigma^{max}_{\operatorname{BO}}(c,\mathcal{X})$. We have m1m-1 local constraints (one for each losing candidate) which can be combined as $\sigma^{min}_{\operatorname{BO}}(w,\mathcal{X})\geq\max_{c\neq w}\sigma^{max}_{\operatorname{BO}}(c,\mathcal{X})$. This transforms into $|\mathcal{X}|\geq\sum_{c}\mu_{\mathcal{X}}(w,c)\geq n(m-1)-\min_{c\neq w}(\mu_{\mathcal{X}}(w,c))$. Finally, since $\mu_{\mathcal{X}}\leq\mu$, $|\mathcal{X}|\geq n(m-1)-\min_{c\neq w}(\mu(w,c))$. ∎
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$ with a Borda score of $\sigma_{w}$. If $\sigma_{w}\leq n(m-1)-\min_{c\neq w}\mu(w,c)$, there exists an MS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $|\mathcal{X}|{=}(m-1)(n(m-1)-\sigma_{w})$.
@@
Suppose $\sigma_{w}\leq n(m-1)-\min_{c\neq w}\mu(w,c)$. Hence, for all $c\in\mathcal{C}\setminus{\{w\}}$, $\delta(c,{E}^{w\rightarrow})=\sigma_{\operatorname{BO}}^{min}(c,G)-\sigma_{\operatorname{BO}}^{max}(c^{\prime},G)=\sigma_{w}-(n(m-1)-\mu(w,c))\leq 0$ and ${E}^{w\rightarrow}$ is at best an MS if all deltas are null else it is a subset of an MS. We start with $\mathcal{X}={E}^{w\rightarrow}$. The idea is then to simply add defeats to candidates that can still achieve more victories than the tournament winner, i.e., for each $c\in\mathcal{C}\setminus{\{w\}}$, add elements of ${E}^{\rightarrow c}$ until $\delta(c,{E}^{w\rightarrow})=0$. In the resulting minimal support $\mathcal{X}$, each losing candidate would have $(n(m-1)-\sigma_{w})$ defeats among which all the tournament winner victories can be found. Thus, $|\mathcal{X}|{=}(m-1)(n(m-1)-\sigma_{w})$. ∎
Before working we the other cases, we show the existence of an SMS which is either included in ${E}^{w\rightarrow}$ or contains it.
@@
Given $n$ voters, an $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$, there exists an SMS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $\mathcal{X}\subseteq{E}^{w\rightarrow}$ or ${E}^{w\rightarrow}\subseteq\mathcal{X}$.
@@
The idea is to go from one SMS to another by adding a win in pairwise comparison for $w$ and removing a lose for another candidate. This can be done as long as there are win for $w$ not in the SMS and defeats for other candidates in the SMS. When one of those two conditions is not met anymore we stop. First we show that: (a) For all SMSs $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $\exists e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\in E\setminus{{E}^{w\rightarrow}}$ and $\exists e_{w}\in{E}^{w\rightarrow}$ with $e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\in\mathcal{X}$ and $e_{w}\not\in\mathcal{X}$, $\left(\mathcal{X}\setminus{\{e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\}}\right)\cup\{e_{w}\}$ is another SMS for $w\in\operatorname{BO}(G)$. Let $\mathcal{X}^{\prime}=\left(\mathcal{X}\setminus{\{e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\}}\right)\cup\{e_{w}\}$. Let $c_{0}\in\mathcal{C}$ s.t. $e_{\mkern 1.5mu\overline{\mkern-1.5mu\!w\mkern-1.5mu}\mkern 1.5mu}\in{E}^{c_{0}\rightarrow}$. We have $\sigma^{min}(c,\mathcal{X}^{\prime})=\sigma^{min}(c,\mathcal{X})+1$, for all $c\in\mathcal{C}\setminus{\{w,c_{0}\}}$, $\sigma^{max}(c,\mathcal{X})=\sigma^{max}(c,\mathcal{X}^{\prime})$ and $\sigma^{max}(c_{0},\mathcal{X})=\sigma^{max}(c,\mathcal{X}^{\prime})+1$. All in all, for all $c\in\mathcal{C}\setminus{\{w\}}$, $\delta(c,\mathcal{X}^{\prime})\geq\delta(c,\mathcal{X})$. Hence, $\mathcal{X}^{\prime}$ is a weak MS for $w\in\operatorname{BO}(G)$. Additionally, $|\mathcal{X}|=|\mathcal{X}^{\prime}|$ and $\mathcal{X}$ is an SMS for $w\in\operatorname{BO}(G)$. Thus, $\mathcal{X}^{\prime}$ is one too.
Finally, given any SMS $\mathcal{X}$ of $w\in\operatorname{BO}(G)$, applying (a) $\min(|\mathcal{X}\setminus{{E}^{w\rightarrow}}|,|{E}^{w\rightarrow}\setminus{\mathcal{X}}|)$ times, we obtain an SMS for $w\in\operatorname{BO}(G)$ such that $\mathcal{X}\subseteq{E}^{w\rightarrow}$ or ${E}^{w\rightarrow}\subseteq\mathcal{X}$. ∎
@@
Let $\mathcal{X}$ be an SMS for $w\in\operatorname{BO}(G)$. Every element of ${E}^{w\rightarrow}$ contributes 1 to every delta (as a victory for the tournament winner, it increases its minimal possible score by 1) and 1 to the specific candidate $c$ it is pointing toward ((w,c) as defeat for c decreases its maximal possible score by 1). Hence, removing an element of ${E}^{w\rightarrow}$ reduces one delta by 2 and the others by 1. By subset minimality its is not possible to remove any extra elements from $\mathcal{X}$ either when there is no delta greater than 2 or a null delta. ∎
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m\geq 2$, and a winning candidate $w\in\operatorname{BO}(G)$ with a Borda score of $\sigma_{w}$. If $\sigma_{w}>n(m-1)-\min_{c\neq w}\mu(w,c)$ and $\forall c\in\mathcal{C}\setminus{\{w\}},\,\delta(c,\mathcal{X})\leq 1$, there exists an MS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $|\mathcal{X}|=\lceil n(m-1)^{2}/m\rceil$.
@@
Lemma 28 gives us the existence of an SMS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $\mathcal{X}\subseteq{E}^{w\rightarrow}$. Suppose $\forall c\in\mathcal{C}\setminus{\{w\}},\,\delta(c,\mathcal{X})\leq 1$.
@@
Observe that the left hand-side of the last equality is an integer and since (m1)/m(m-1)/m is strictly less than 1 and non zero ($m\geq 2$), $\sum_{c\neq w}\mu_{\mathcal{X}}(w,c)\leq n(m-1)^{2}/m+(m-1)/m\implies|\mathcal{X}|=\sum_{c\neq w}\mu_{\mathcal{X}}(w,c)\leq\lceil n(m-1)^{2}/m\rceil$. ∎
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$ with a Borda score of $\sigma_{w}$. If $\sigma_{w}>n(m-1)-\min_{c\neq w}\mu(w,c)$ and $\exists c_{0}\in\mathcal{C}\setminus{\{w\}},\,\delta(c_{0},\mathcal{X})\leq 0$, there exists an MS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $|\mathcal{X}|\leq n(m-1)-\min_{c\neq w}\mu(w,c)$.
@@
Lemma 28 gives us the existence of an SMS $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ such that $\mathcal{X}\subseteq{E}^{w\rightarrow}$. Suppose $\exists c_{0}\in\mathcal{C}\setminus{\{w\}},\,\delta(c_{0},\mathcal{X})\leq 0$. Since $\mathcal{X}$ is an MS, $\forall c\in\mathcal{C}\setminus{\{w\}},\,\delta(c,\mathcal{X})\geq 0$. Hence, $\delta(c_{0},\mathcal{X})=0$, i.e., $\sum_{c\in\mathcal{C}\setminus{\{w\}}}\mu_{\mathcal{X}}(w,c)=n(m-1)-\mu_{\mathcal{X}}(w,c_{0})$. Additionally, $\forall c\in\mathcal{C}\setminus{\{w\}},\,\delta(c,\mathcal{X})\geq 0=\delta(c_{0},\mathcal{X})$ and we have $\forall c\in\mathcal{C}\setminus{\{w\}},\,\mu_{\mathcal{X}}(w,c)\geq\mu_{\mathcal{X}}(w,c_{0})$. Hence, we have $\forall c\in\mathcal{C}\setminus{\{w\}},\,|\mathcal{X}|\leq n(m-1)-\mu_{\mathcal{X}}(w,c)$ and in particular, $|\mathcal{X}|\leq n(m-1)-\min_{c\in\mathcal{C}\setminus{\{w\}}}\mu_{\mathcal{X}}(w,c)$.
@@
Note that in the last formula we have $\min_{c\in\mathcal{C}\setminus{\{w\}}}\mu_{\mathcal{X}}(w,c)$ and not $\min_{c\in\mathcal{C}\setminus{\{w\}}}\mu(w,c)$. We still have to show that there is an SMS such that there exists a candidate $c$ distinct from $w$ such that $\mu_{\mathcal{X}}(w,c)=\mu(w,c)$.
@@
Suppose for all SMSs $\mathcal{Y}\subseteq{E}^{w\rightarrow}$, $|\mathcal{Y}|>n(m-1)-\min(\lfloor n(1-1/m)\rfloor,\min_{c\neq w}\mu(w,c))$. If $\forall c\in\mathcal{C}\setminus{\{w\}}$, $\delta(c,\mathcal{Y})\leq 1$, we are done thanks to Lemma 30. Otherwise, we have the existence of $c_{1}\neq w$ such that $\delta(c_{1},\mathcal{Y})>1$ and by minimality of MSs, $\exists c_{0}\neq w$ such that $\delta(c_{0},\mathcal{Y})=0$. Suppose now that $\forall c\neq w$ such that $\delta(c,w)=0$, $\mu_{\mathcal{Y}}(w,c)<\mu(w,c)$. Consider $\mathcal{Y}^{\prime}$ such that $\mu_{\mathcal{Y}^{\prime}}(w,c_{0})=\mu_{\mathcal{Y}}(w,c_{0})+1$, $\mu_{\mathcal{Y}^{\prime}}(w,c_{1})=\mu_{\mathcal{Y}}(w,c_{1})-1$ (which is possible as $\mu_{\mathcal{Y}^{\prime}}(w,c_{1})<\mu(w,c_{1})$) and else $\mu_{\mathcal{Y}^{\prime}}=\mu_{\mathcal{Y}}$. $\delta(c_{0},\mathcal{Y}^{\prime})=1$, $\delta(c_{1},\mathcal{Y}^{\prime})=\delta(c_{1},\mathcal{Y})-1\geq 1$ and $\forall c\not\in\{w,c_{0},c_{1}\},\,\delta(c,\mathcal{Y}^{\prime})=\delta(c_{1},\mathcal{Y})\geq 1$. Since all deltas are positive, $\mathcal{Y}^{\prime}$ is a weak minimal support and since $|\mathcal{Y}^{\prime}|=|\mathcal{Y}|$ and $\mathcal{Y}$ is an SMS, $\mathcal{Y}^{\prime}$ is an SMS too. By passing from $\mathcal{Y}$ to $\mathcal{Y}^{\prime}$ we strictly reduce the number of losing candidates with a null delta ($\delta(c_{0},\mathcal{Y})=0$ but $\delta(c_{0},\mathcal{Y}^{\prime})=1$). Since the number of candidates is finite, this process has to stop at one point, i.e., $\exists\mathcal{Y}^{*}$ an SMS such that $\exists c^{*}\neq w$ with $\delta(c^{*},\mathcal{Y}^{*})=0$ and $\mu_{\mathcal{Y}}(w,c^{*})=\mu(w,c^{*})$. We have $\sum_{c\neq w}\mu_{\mathcal{Y}^{*}}(w,c^{*})-n(m-1)+\mu(w,c^{*})=0$ and $|\mathcal{Y}^{*}|=n(m-1)-\mu(w,c^{*})$. Finally, since $\min_{c\neq w}\mu(w,c)\leq\mu(w,c^{*})$, exists an SMS $\mathcal{X}$ such that, $|\mathcal{X}|\leq n(m-1)-\min_{c\neq w}\mu(w,c)$. ∎